An empty glass bulb has a volume of 350mL and a mass of 82.561 g. When filled with an unknown gas at 733mmHg and 22degrees Celsius, the bulb has a mass of 82.786g. Is the gas argon(Ar), methane(CH4), or nitrogen monoxide(NO)? How do I know which gas it is? How do I find that out?
1
Expert's answer
2019-11-07T08:04:26-0500

V=0.350L

∆m=82.786g-82.561g=0.225g the mass of an added gas

P=733mmHg/760=0.964atm

T=22°C+273.15=295K

Rearrange the ideal gas law to solve for n:

PV=nRT

n=PV/(RT)=(0.964atm*0.350L)/(0.08206atm*L/(mol*K)*295K)=0.0139mol

Find the ratio of the mass to the number of moles of an unknown gas to find its molar mass: 0.225g/0.0139mol=16.1g/mol

The molar mass of Ar is 39.95 g/mol, of NO is 30.01 g/mol, and of CH4 is 16.04 g/mol

Thus, the gas in the bulb is methane(CH4)